# General-Purpose Multicore Architectures

### v1

@@ 7.3 Chiplet-Based Multicore Design
Beyond the cores themselves, the success of SoCs have demonstrated the benefits of maximizing energy efficiency through directed specialization. However, there remains a tension between high degrees of specialization and non-recurring engineering (NRE) costs, such as those involved with design, layout, and verification. As a simple motivating example, let’s revisit the tiled multicore design from Section 3.1. While tiling helps reduce NRE costs, the tile is a fixed design: for one core, we have a fixed amount of L1 instruction and data caches, L2 cache, and potentially LLC slice. If a manufacturer wants to adapt this tile for a platform whose workloads do not exhibit significant locality, they may want to significantly reduce the cache sizes, but doing so requires a new tile to be designed and verified. Moreover, the die that is etched for a multicore CPU will have a fixed number of tiles laid out, again restricting the flexibility of the CPU and requiring non-trivial NRE costs if a die with a different core/tile count is needed.
@@
Chiplets offer four advantages for manufacturing. First, as we already discussed, chiplets allow for modular components that can be reused and resized after the dies have been fabricated, at low cost. Second, chiplets can allow for dies fabricated using different manufacturing process technologies to be connected together into a single package (this is known as heterogeneous integration). If, for example, the cache does not need to be manufactured using the state-of-the-art manufacturing process, manufacturers can reduce costs by fabbing the chip using an older, cheaper process. Third, overall yield increases, because a silicon fault is now isolated to a much smaller chiplet, which can be replaced at much lower cost than disposing of an entire die. Fourth, with the breakdown of Dennard scaling [13, 14], dies are now growing in size to continue scaling up the total transistor count, but these sizes are approaching the reticle limits (i.e., the largest possible chip that can be etched) of our lithography equipment. Chiplets can overcome these limits by allowing for multiple larger chiplet dies to be composed into a package, where the total area of the chiplets is significantly larger than what any one die could be.
@@ 8 Conclusion
The rise of multicore CPUs highlighted a key shift in the computer architecture community, as concerns about thermal dissipation and limitations of ILP hastened a collective change in mindset about the importance of power and the potential of parallel processing. Innovations in multicore CPU design have led to reduced design and verification effort, support for specialized hardware accelerators, and increased composability of modular components. Today, multicore CPUs can be found across most modern computers, ranging from embedded platforms and smartphones, through personal laptops and desktops, to large-scale distributed computing environments. Combined with the emergence of simplified parallel programming frameworks, multicore CPUs have led to commonplace exploitation of thread-level parallelism, delivering significant performance improvements while maintaining reasonable power and energy budgets. We expect multicore CPUs to continue to evolve over the next several decades, as recent trends towards CPU specialization (particularly in the modern CPU landscape with systems-on-chip and chiplet-based fabrication) open up new opportunities for maximizing the efficiency and performance of the next generation of computing platforms.

### v2

@@ 7.3 Chiplet-Based Multicore Design
Beyond the cores themselves, the success of SoCs have demonstrated the benefits of maximizing energy efficiency through directed specialization. However, there remains a tension between high degrees of specialization and non-recurring engineering (NRE) costs, such as those involved with design, layout, and verification. As a simple motivating example, let’s revisit the tiled multicore design from Section 3.1. While tiling helps reduce NRE costs, the tile is a fixed design: for one core, there is a fixed amount of L1 instruction and data caches, L2 cache, and potentially LLC slice. If a manufacturer wants to adapt this tile for a platform whose workloads do not exhibit significant locality, they may want to significantly reduce the cache sizes, but doing so requires a new tile to be designed and verified. Moreover, the die that is etched for a multicore CPU will have a fixed number of tiles laid out, again restricting the flexibility of the CPU and requiring non-trivial NRE costs if a die with a different core/tile count is needed.
@@
Chiplets offer four advantages for manufacturing. First, as already discussed, chiplets allow for modular components that can be reused and resized after the dies have been fabricated, at low cost. Second, chiplets can allow for dies fabricated using different manufacturing process technologies to be connected together into a single package (this is known as heterogeneous integration). If, for example, the cache does not need to be manufactured using the state-of-the-art manufacturing process, manufacturers can reduce costs by fabbing the chip using an older, cheaper process. Third, overall yield increases, because a silicon fault is now isolated to a much smaller chiplet, which can be replaced at much lower cost than disposing of an entire die. Fourth, with the breakdown of Dennard scaling [13, 14], dies are now growing in size to continue scaling up the total transistor count, but these sizes are approaching the reticle limits (i.e., the largest possible chip that can be etched) of our lithography equipment. Chiplets can overcome these limits by allowing for multiple larger chiplet dies to be composed into a package, where the total area of the chiplets is significantly larger than what any one die could be.
@@ 8 Conclusion
The rise of multicore CPUs highlighted a key shift in the computer architecture community, as concerns about thermal dissipation and limitations of ILP hastened a collective change in mindset about the importance of power and the potential of parallel processing. Innovations in multicore CPU design have led to reduced design and verification effort, support for specialized hardware accelerators, and increased composability of modular components. Today, multicore CPUs can be found across most modern computers, ranging from embedded platforms and smartphones, through personal laptops and desktops, to large-scale distributed computing environments. Combined with the emergence of simplified parallel programming frameworks, multicore CPUs have led to commonplace exploitation of thread-level parallelism, delivering significant performance improvements while maintaining reasonable power and energy budgets. Multicore CPUs are expected to continue evolving over the next several decades, as recent trends towards CPU specialization (particularly in the modern CPU landscape with systems-on-chip and chiplet-based fabrication) open up new opportunities for maximizing the efficiency and performance of the next generation of computing platforms.